

********************************************************************
UARS WEEKLY STATUS REPORT - 8 MAY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and
occultation scan data without apparent problems.
***********
PEM
***********
PEM is not active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is under evaluation by PEM.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Apr 29, in the reduced 85 watt power
mode. MLS instrument operations have continued flawlessly. MLS was powered
off May 2 at 23:32 UT. MLS is scheduled to begin south viewing operations
starting May 20.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF receieved Level 0 for UARS days 2417, 2418, 2419 and 2420 on 05/03 and
05/04. These days were delayed due to the building power outage on 04/24
through 04/27.
The CDHF has received a total of 386 CD-ROM requests as of 08-May-1998.
No new orders were received.
The CDHF receieved UKMO Day 2263 and 2264.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
43 days PEM (3.32) PROD (1379R-1212C=167L)
2 days HAL (4.80) PROD (Test) (2R-2C=0L)
1 days WIN (5.1) ENG (1R-1C=0L)
CDHF worked on proposed UARS CDHF system configuration changes and
associated improvements and impacts.
***********
NOAA DATA
***********
NOTICE OF CHANGE IN TEMPERATURES FROM NCEP ANALYSES
On May 1, 1998 NCEP operational stratospheric analyses began using RTOVS
(Revised TIROS Operational Vertical Soundings) from the NOAA-14 satellite.
TOVS data had been produced by NESDIS (National Environmental Satellite Data
and Information Service) from the operational NOAA series of satellites for the
last two decades. RTOVS was introduced as a transition to ATOVS (Advanced
TOVS), in preparation for the launch of NOAA-K later this month (May 13, 1998).
Once operational, NOAA-K becomes NOAA-15. TOVS and RTOVS soundings use data
from the Stratospheric Sounding Unit (SSU), the Microwave Sounding Unit (MSU),
and the High resolution InfraRed Sounder (HIRS). ATOVS will derive soundings
from the HIRS and the Advanced Microwave Sounding Unit (AMSU). Comparisons to
radiosonde data show general improvement with RTOVS data. Average RTOVS/TOVS
differences at the various pressure levels change in magnitude and sign with
season and latitude, but in general are less than 1 K degree. In some
instances, differences of up to 3 K apparent.
For more information, contact Mel Gelman at melvyn.gelman@noaa.gov or
at 301-763-8000 x7558.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 April 1998 (Orbit 36169, GMT Day 115) through 01 May
1998 (Orbit 36272, GMT Day 121).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 115/0205 36170 On
HALOE 118/2304 36228 Off
MLS 119/0039 36229 On
HRDI 120/1115 36250 Off
A summary of Instrument operations supplied by the Mission Planning Group
for the month of April 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle is now decreasing toward a minimum angle of 0.0 degrees on
05 May 1998 (DOY 125).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.2 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.2 degrees C. Additional battery performance data for
each battery on 25 April 1998 and 01 May 1998 are provided in Appendix D.
The clock error ranged from -10.9 to +11.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 08
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HALOE 01
% HRDI 02
% MLS 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/062 - 2287.496945
Transponder B: 98/062 - 2287.497010
GROUND SYSTEM OPERATION
No (0) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/120 show the observatory data loss to be
20 hours, 43 minutes, 08.788 seconds (an increase of 14.624 seconds since
last report period). This is a 0.0358 percent data loss which equals a
99.9642 percent data capture for the mission.
The increase of 14.624 seconds occurred over a three day period:
7 9.888 seconds on Day 115 due to hits in the data,
7 0.448 seconds on Day 116 due to hits in the data, and
7 4.288 seconds on Day 117 due to hits in the data.
The FOT EXPERIENCED along-track ephemeris transition comparisons in excess
of 1.5 km when processing the ephemeris in the Command Management System
(CMS) software this week. The out-of-limit comparisons occurred on two of
the three ephemeris delivery days. The FOT will contact FDF next week to
discuss the status of their investigation into the need for more frequent
spacecraft ephemeris deliveries.
The FOT performed a limited amount of IP testing this week. As with
previous testing, the FOT ran a spare AP in a parallel environment to
monitor I-channel return telemetry. The FOT also attempted to monitor Q-
channel OBC dumps via the IP system, but was unsuccessful. On one event,
the TAC crashed before the Q-channel dump was started. On a second event,
the TAC locked on the Q-channel data, but the dump compare was never
conducted on the AP. This was probably due to an operational or AP
configuration problem by the FOT. Investigation is underway as to what the
FOT must due to achieve a successful Q-channel OBC dump. Further testing
is planned for next week
The FOT continued making modifications to the OBCSTF resident POCC database
this week. The number of additions, deletions, and modifications has
turned out to be far more extensive than originally thought. In addition,
the FOT has been having intermittent problems with the OBCSTF terminal
Rlocking upS during database entry. The majority of the database
modifications have been completed, but many hours worth of work still
remain. The database work is expected to be completed by 06 May 1998.
The FOT made a few modifications to the new CMS versions of Relative Time
Sequences (RTSs) 39 and 40 this week. The changes were necessary to
preclude the need for frequent RTS reloading to properly control the
SOLSTICE instrument. The modifications were coordinated the SOLSTICE
operations team.
The FOT received requests from SOLSTICE operations to perform two manually
generated SSPP Stellar Offset experiments on GMT days 125 and 127.
The second Building 23 power outage began on schedule at noon on 24 April
1998 and continued without event until 27 April 1998. All data has been
recovered from White Sands Ground Terminal.
Four days of level-0 data is missing (24 April to 27 April) on the CDHF
system due to the weekend power outage at the close of this report period.
RDummyS telemetry files were created for the missing days to allow daily
UGPL products to run during the week. The missing days are planned to be
processed over the weekend of 02-03 May, and the RdummyS files will be
replaced with actual data the following Monday (04 May). Post-status note:
All missing level-0 data was processed by about noon today (04 May), and
UGPL products have been restored to nominal operation. All data is
included in the data quality statistics at the beginning of this section.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) continued. FOT
coordinated with PEM, MLS, and SOLSTICE for inputs to new PMON software
(these instruments primarily affected by changes) this week. New
operations scenarios for each of these instruments, including new ground
database commands, telemetry, and RTS/PDB definitions were generated.
Work continues on the UARS August orbit adjust planning.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
A response to a list of questions was prepared for distribution this week.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
A follow-up meeting to discuss and confirm the changes made to the OBC
Ground Image to support the OBC Flight Software update was held on 30 April
1998. A decision was reached to proceed with the development/release of
Version 10.11 of the Application Processor software to support the OBC PMON
update. An operational pack will be available by the end of May 1998. FOT
will conduct testing of the software and data base changes using the 10.11
version test pack while the official release process and operational pack
generation takes place.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 05 May 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 May 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
36170 115/0205 98-149 HRDI Power ON
36180 115/1736 98-150 OBC Clk Adj fr 0.00 to 20.6 msec/day
36209 117/1636 98-151 OBC Clk Adj fr 20.6 to 0.00 msec/day
36228 118/2304 98-152 HALOE Power OFF
36229 119/0039 98-153 MLS Power ON
36250 120/1115 98-154 HRDI HRDI to low power
36272 121/2136 98-155 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 May 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI PTS Pulse Bus Current Monitor 36171,235,39
QTSISAI/F Y-HI SSPP Baseplate by SISA I/F 36176,89,203,204,
267-72
WIEUTEMP Y-LO WINDII EU Temperature 36220,22,26,27,29-35
NBRECBI Y-HI NBTR B Total Current 36224,39
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 36224,39,49,69,70
CDUNRG28 Y-LO +28V Unregulated Bus 36249
HRIFESURVTMP Y-HI IFE Survival Heater Temp 36253
HRRADTEMP Y-HI IFE Radiator Temperature 36253
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 May 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36180 115/17:42 (25 Apr 98) from 00.0 to 20.6 msec/day
36209 117/16:50 (27 Apr 98) from 20.6 to 00.0 msec/day
36272 121/21:46 (01 May 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 May 1998
25 Apr 1998 - GMT Day 115 Beta = 36.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +56.0/ -11.2 +6.44/+4.51 +29.6/-16.0 25.0 1.022 22.4
3 +123.0/ +0.0 +5.15/+3.23 +28.8/-16.0 24.8 1.024 21.9
01 May 1998 - GMT Day 121 Beta = 17.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.73/-3.99 +0.0/ +0.0 18.4 0.000 0.0
2 +39.2/ -5.6 +7.09/+5.47 +32.0/-15.6 24.2 1.019 23.8
3 +44.8/ -5.6 +5.79/+4.51 +30.4/-16.0 24.2 1.021 23.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 May 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-022 FOT 36272 121/2136 N/A TAC #2 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-015 FOT April 1998 14170 GENERIC LATE ACQ
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 Generic Late Acq for May 1998 (Open until 31 May 1998).
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 01 May 1998
INSTRUMENT SUMMARY FOR MONTH OF April 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- ------------
ACRIM 598 122 0 0
HALOE 500 220 0 0
HRDI 391 316 13 0
MLS 236 484 0 0
PEM (axis) 0 720 0 0
PEM (zeps) 111 609 0 0
PEM (neps) 0 720 0 0
SOLSTICE 720 0 0 0 Note 1
SUSIM 587 133 0 0
WINDII 185 535 0 0
Note 1 - No star data March 15-19
Cold side of spacecraft was facing northward: 28.5 days
Cold side of spacecraft was facing southward: 1.5 days
SEVNTFW18